How can I determine the LCD by inspection and by division of prime numbers, such as 3/4,7/12,5/6,1/5?
Thank you very much.
Ellie

Answer by stanbon(75887)

How can I determine the LCD by inspection and by division of prime numbers, such as 3/4,7/12,5/6,1/5?
-----
Prime factor form of the denominators:
4 = 2^2
12 = 2^2*3
6 = 2*3
5 = 5
------
The LCD must contain each prime factor in its highest power:
LCD = 2^2*3*5 = 60
-----
Convert each fraction to a fraction with the LCD as its denominator:
3/4 = (3/4)(15/15) = 12/60
7/12 = (7/12)(5/5) = 35/60
5/6 = (5/6)(10/10) = 50/60
1/5 = (1/5)(12/12) = 12/60
===============================
cheers,
Stan H.